OpenStudy (anonymous):

help with finding total coefficient for x^0 for binomial expansion of (X^2+1/x)^12.. i get the coefficient is 12C0 x^2^0 ((((or 1)))) * 1/x^12... but my book says 495...how do i get that?

OpenStudy (campbell_st):

its when \[(x^2)^{12-r}(x^{-1})^r = x^0\] equating the powers 24 - 2r - r = 0 24 = 3r r = 8 so then you need to find \[^{12}C _{8} (x^2)^{12 - 8} (x^{-1})^8\] gives \[495(x^8)(x^{-8}) = 495\]

OpenStudy (anonymous):

thank you, but why if its the coefficient of x^0 term, why is it not just the final term?

OpenStudy (anonymous):

or does it mean when simply all X's will cancel out from the binomial

OpenStudy (campbell_st):

the final term is \[^{12}C _{12}(x^2)^0(x^{-1})^12 = ^{12}C _{12}x^{-12}\]

OpenStudy (campbell_st):

its when the product of the terms results in a power of zero... remember multiplying the same pronumeral.... add the powers...

OpenStudy (anonymous):

so if i wanted when x^4, just set it equal to 4. awesome thanks so much

OpenStudy (campbell_st):

yep it would be 24 - 2r - r = 4 or 24 - 3r = 4 but it you can't get 4 the powers will be 24, 21, 18, 15, ... to -12 etc

OpenStudy (anonymous):

This is the whole expansion \[ \left(x^2+\frac{1}{x}\right)^{12}= x^{24}+12 x^{21}+66 x^{18}+220 x^{15}+\\495 x^{12}+\frac{1}{x^{12}}+\\792 x^9+\frac{12}{x^9}+924 x^6+\frac{66}{x^6}+792 x^3+\\\frac{220}{x^3}+495 \]

OpenStudy (anonymous):

The method described by @campbell_st is the way to do it.